# Self-Modeling Robots by Photographing

### v1

@@ 4.1 Simulation
We also present results without joint limits when posing the robot, as shown in Figure 4b. For each joint configuration, we display two images from different random view angles, with at least one of the robot arm’s seven joints exceeding $\frac{\pi}{2}$. Self-modeling with a larger range of joint angles is significantly more challenging, leading to results that are worse than those obtained within a smaller range. The robot’s end effector is not precisely reconstructed due to its position at the end of the kinematic chain and the complexity of its movement. However, other parts of the robot are still accurately modeled. Compared to the method proposed by Schulze and Lipson (2024), which uses a NeRF-based method for self-modeling, our method reconstructs the robot’s shape more precisely, allowing easier identification of the robot’s texture and morphology in the rendered images.
We also present results without joint limits when posing the robot, as shown in Figure 4b. For each joint configuration, we display two images from different random view angles, with at least one of the robot arm’s seven joints exceeding $\frac{\pi}{2}$. Self-modeling with a larger range of joint angles is significantly more challenging, leading to results that are worse than those obtained within a smaller range. The robot’s end effector is not precisely reconstructed due to its position at the end of the kinematic chain and the complexity of its movement. However, other parts of the robot are still accurately modeled. Compared to the results presented by Schulze and Lipson (2024), which uses a NeRF-based method for self-modeling, our method reconstructs the robot’s shape more precisely, allowing easier identification of the robot’s texture and morphology in the rendered images.
@@ 4.2 Physical experiment
We utilized a smartphone and an Omnivision OV5640 color sensor attached to the end effector of a robot to capture images of the OpenManipulator robot. All images were resized to a resolution of $640\times 480$ while maintaining the aspect ratio. Following a procedure similar to that in the simulation, we first used both cameras to capture images with the robot in its zero pose. The smartphone was free to capture images from various view angles, whereas the color sensor could only capture images from fixed angles. To simplify data collection, we captured images in a forward-facing manner instead of a $360^{\circ}$ inward approach. Notably, multiple fixed cameras could be used in place of the robot’s color sensor during data collection. Camera parameters were computed using structure-from-motion (SfM) (Schönberger and Frahm, 2016) based on the static scene. Subsequently, we randomly posed the robot and used only the color sensor to capture images from the calculated view angles. We restricted the joint angles between $[-\frac{\pi}{6},\frac{\pi}{6}]$ to capture 100 images, then posed the robot without joint limitations to capture an additional 1,000 images. We employed the segment anything model (SAM) (Kirillov et al., 2023) to generate masks and segment the robot from the background.
@@
The results of physical experiments are shown in Figure 5. We randomly selected three sets of joint angles that were not part of the training data, rendered the corresponding robot images using the trained model, and compared them to the ground truth. The comparison reveals that the proposed method accurately represents the robot’s morphology and texture, further validating the effectiveness of our approach.
Conducting self-modeling with real-world data is more challenging than with simulation data. Thus, the real-world results are not as accurate as those obtained from simulations. This can be attributed to several factors: (1) Potential inaccuracies in the camera parameters calculated by SfM, (2) Imperfect segmentation of the robots from the background in certain images, and (3) A smaller volume of data is available in real-world experiments compared to simulation. Nevertheless, our real-world results are significantly closer to the ground truth compared to previous work by Chen et al. (2022).
@@ 4.3 Applications
In this section, we use the proposed self-model to complete some downstream tasks: reaching a target and reaching a target while avoiding an obstacle. We also show that our self-model can solve the inverse kinematics problem.
@@ 4.3.1 Reaching a target
A green ball was placed within the robot’s workspace, requiring the robot to use its self-model to determine how to make contact with the ball using a specific part of its body. Our neural bone representation simplifies identifying specific parts of the robot and enables targeted control to reach the desired object. We use the robot’s end effector as an example. Starting from a resting configuration, the joint angles are progressively adjusted to minimize the distance between the target and the center of the neural bone associated with the end effector. This is accomplished by fixing the self-model parameters and applying gradient descent to automatically optimize the joint angles. In practice, the AdamW (Loshchilov, 2017) optimizer is utilized, and the loss function is defined as the mean squared error (MSE) between the bone center and the target position. Unlike method of Chen et al. (2022), by altering the neural bone applied in the loss function, the robot can be controlled to reach the target using different parts without the need for additional training.
@@ 4.3.2 Reaching a target while avoiding an obstacle
An obstacle was placed at a location the robot would pass through after completing the first task of reaching the target. The robot is required to find a new joint configuration to reach the target, as the original configuration would result in a collision with the obstacle. To address this, an additional loss function is introduced alongside the MSE loss function used in the previous task. All 3D Gaussians are treated as point clouds, and the distances $D$ between the points and the obstacle are computed. The loss function is then formulated as:
@@
where $\varepsilon$ is a threshold that controls how close the robot can get to the obstacle. In our experiment, about 50,000 3D Gaussians were used in the self-model. All 3D Gaussians were forced to keep a distance from the obstacle.
The two applications above are demonstrated in Figure 6. Images in the first row show how our self-model controls the end effector of the robot to reach the target, shown here as a green ball. The robot’s final pose in this task is shown in the fifth image of the first row (Movie S1). Next, an obstacle was placed at the position of the robot’s fourth link in this final pose, and images in the second and third rows show two perspectives of our self-model guiding the robot to complete the target-reaching task while avoiding the obstacle. Clearly, the robot adopts a new pose to avoid the obstacle and reach the target (Movie S2 and Movie S3). The minimum distance between the obstacle and the robot’s 3D Gaussian representations is also shown in the last row of Figure 6b.
Figure 6: Results of applications. a Snapshots and joint angles of the robot with respect to step number. b Snapshots from different perspectives, joint angles of the robot, and the minimum distance between the obstacle and the robot with respect to step number. The unit of the joint angle is radian. At the beginning, all joints of the robot are set to 0. Using solely the self-model, the robot is able to find joint configurations that can fulfill the tasks. c Results of inverse kinematics estimation. The ground truth joints are [-0.204, -0.9319, -0.3448, -1.5012, 0.4474, 0.105, 0.3166]. The estimated joints are [-0.2127, -0.9321, -0.3051, -1.5050, 0.4863, 0.0258, 0.3113]. Joint angles in radians are slightly different, but the morphology of the robot is close to the ground truth.
@@ 4.3.3 Robot inverse kinematics estimation
With our self-model, we can estimate the robot’s inverse kinematics. Given images of the robot, our self-model can predict a joint configuration to pose the robot similarly. After freezing self-model parameters and treating joint angles as optimizable, we employ gradient descent to approximate the joint configuration. Comparing rendered and ground truth images, we optimize joint angles to achieve reasonable values. In practice, we use two images to estimate the robot’s pose. Our estimation results are illustrated in Figure 6c. Although estimated joint angles differ slightly from the actual angles, rendered images at the estimated joint angles match ground truth images. Our self-model effectively poses the robot in alignment with the ground truth. Additional images from varied view angles could further improve joint angle estimation accuracy.
@@ 5 Discussion
We propose a 3DGS-based robot self-modeling method that eliminates the need for depth information. This approach achieves high-quality modeling of the robot’s morphology, texture, and kinematics. Our method initiates modeling from the robot’s rest pose. We employ learnable neural ellipsoid bones of similar volume to represent morphology and texture. A kinematic network is trained to capture the robot’s kinematic properties. Each 3D Gaussian’s movement is governed by transformations of neural bones, enabling full model control. Rendering 3D Gaussians into images enables us to visualize the robot’s morphology and texture. Our self-model accurately reconstructs the robot across various joint configurations, achieving high-quality image synthesis.
The self-model developed using our method can be applied in downstream tasks. With this self-model, the robot is capable of solving motion-planning tasks. Our neural bone representation directly models the robot’s structure, allowing us to control different parts of the robot by specifying which bone to optimize in the loss function without requiring additional information or training to adjust the self-model. Since our self-model can synthesize high-quality images, it can also estimate the robot’s inverse kinematics by replicating input image poses.
@@
However, our method can be further improved in several ways. (1) A better initialization of our neural bones is possible. The current method requires a predefined number of bones to initialize the structure. We also assume that each bone is similar in size, but in reality, the size of robot links may vary, necessitating multiple bones to represent a single link. (2) Our method may not always accurately reconstruct the robot at certain joint configurations, especially when the joint’s movement range is unrestricted. It is worth researching how to obtain a more accurate self-model across all joint configurations. (3) It is important to find a method to map the 3DGS self-model to a mesh model and modify it for compatibility with common robot control pipelines, such as MOVEIT (Coleman et al., 2014) or current reinforcement learning toolkits. (4) We need multiple views to reconstruct the model, and the number of views, along with the camera parameters, can affect the final result. Researching how to self-model the robot with high quality using fewer views and without known camera parameters is a valuable direction to explore.
6 Author Contributions
7 Statements and declarations
7.1 Ethical considerations
Not applicable.
7.2 Consent to participate
Not applicable.
7.3 Consent for publication
@@
Not applicable. {dci} The author(s) declared no potential conflicts of interest with respect to the research, authorship, and/or publication of this article.
@@
Movie S2: The robot is controlled by its self-model to reach the green target while avoiding the red obstacle at viewpoint 1.
Movie S3: The robot is controlled by its self-model to reach the green target while avoiding the red obstacle at viewpoint 2.

### v2

@@ 4.1 Simulation
[…]
The quantitative results of the above four different joint configurations are presented in Table 1. The metrics we used are peak signal-to-noise ratio (PSNR) (Hore and Ziou, 2010), structural similarity (SSIM) (Odena et al., 2017; Wang et al., 2004), and learned perceptual image patch similarity (LPIPS) (Zhang et al., 2018), which are commonly used to evaluate the quality of synthetic images. The quantitative results for each joint configuration were calculated using seven images, consisting of six horizontal views around the robot and one top view. We also validated the proposed method in a test set containing 2,000 images captured from random view angles around the robot, with varying joint configurations. The results of the test set indicate that our self-model can correctly synthesize images of the robot at different joint configurations and view angles. It is worth noting that larger changes in joint configurations make it more challenging to obtain an accurate self-model, and the test set includes joint configurations with minor changes. Consequently, the results for the 4 example joint configurations are worse than those of the test set. We also computed the Chamfer distance between the extracted mesh and the ground truth mesh for the four demonstrated joint configurations. We uniformly sampled 10,000 points on the mesh, and used the PyTorch3D (Johnson et al., 2020) library to calculate the Chamfer distance. The calculation of Chamfer distance can be formulated as:
$CD(S_{1},S_{2})$ $=\frac{1}{|S_{1}|}\sum_{x\in S_{1}}\min_{y\in S_{2}}\left\|x-y\right\|$ (25)
$+\frac{1}{|S_{2}|}\sum_{y\in S_{2}}\min_{x\in S_{1}}\left\|y-x\right\|$
where $S_{1}$ and $S_{2}$ are two point clouds sampled from two meshes.
As noted previously, our method does not provide a direct 3D representation. Consequently, meshes extracted from the model may not accurately reflect its modeling quality, which reduces the utility of using the Chamfer distance metric for evaluation.
We also present results without joint limits when posing the robot, as shown in Figure 6(b). For each joint configuration, we display two images from different random view angles, with at least one of the joint angles exceeding $\frac{\pi}{2}$. Self-modeling with a larger range of joint angles is significantly more challenging, leading to results that are worse than those obtained within a smaller range. The robot’s end effector is not precisely reconstructed due to its position at the end of the kinematic chain and the complexity of its movement. However, other parts of the robot are still accurately modeled. Furthermore, we conducted a comparison of the meshes generated by our method, the method proposed by Chen et al. (2022), and the method proposed by Schulze and Lipson (2024). These meshes are illustrated in Figure 7. We also calculated the Chamfer distance as listed in Table 2. Note that the chamfer distance of the method proposed by Schulze and Lipson (2024) is directly taken from their paper. We extracted meshes of the robot at the same joint configurations presented by Schulze and Lipson (2024), and we used the same annotations. However, $\theta^{(b)}$ was omitted for the sake of rigor, because a significant discrepancy between our model’s results and the provided ground truth is observed at only this joint configuration and we assumed the value they provided is incorrect. After changing the second angle value from a negative to a positive value (from [0.89,1.26,2.36,0.62,1.92,2.47,1.77][-0.89,1.26,2.36,-0.62,-1.92,2.47,-1.77] to [0.89,1.26,2.36,0.62,1.92,2.47,1.77][-0.89,-1.26,2.36,-0.62,-1.92,2.47,-1.77]), the results exhibited closer agreement with the ground truth. We illustrated this in Figure 8. We implemented the method in Chen et al. (2022) based on the code released by them to generate meshes and adapted it to the 7-DOF robot used in our simulation by simply changing the input dimension to 7. The training took about 2 days on a single Nvidia RTX 3090 GPU. As for the method in Schulze and Lipson (2024), since the robot and joint angles we used are the same, we directly compared the results in their paper. Compared to the method proposed by Schulze and Lipson (2024), which uses a NeRF-based method for self-modeling, and the method proposed by Chen et al. (2022), which uses an SDF-based method and requires depth information, our method reconstructs the robot’s shape more precisely, allowing easier identification of the robot’s surface color and morphology. Nevertheless, one may notice that in configuration $\theta^{(c)}$, the method of Chen et al. (2022) attains the lowest chamfer distance. A possible reason is that as we sample 10,000 points on the mesh, the mesh quality for the links near the robot’s base is very high, and they account for a large portion of the model. However, near the end of the robot, the disconnected section has a ground truth mesh that is inherently thin, resulting in fewer collected data points for that area. Consequently, when the metric is averaged, this leads to a lower overall Chamfer distance.
Figure 7: Mesh comparison of our method, the method in Chen et al. (2022), and the method in Schulze and Lipson (2024). Meshes provided by our method are more easily identifiable by human perception.
Figure 8: a Ground truth we rendered and meshes extracted under the changed joint configuration. b Images of ground truth and extracted mesh provided by Schulze and Lipson (2024). c Ground truth we rendered and mesh extracted under the original provided joint configuration.
@@ 4.2 Physical experiment
We utilized a smartphone and an Omnivision OV5640 color sensor attached to the end effector of a robot to capture images of the OpenManipulator robot. All images were resized to a resolution of $640\times 480$ while maintaining the aspect ratio. Following a procedure similar to that in the simulation, we first used both cameras to capture images with the robot in its zero pose. The smartphone was free to capture images from various view angles, whereas the color sensor could only capture images from fixed angles. To simplify data collection, we captured images in a forward-facing manner instead of a $360^{\circ}$ inward approach, which means we used multiple fixed cameras positioned within a certain angular range in front of the robot to capture the images of the robot. Notably, multiple fixed cameras could be used to accelerate the data collection process. Camera parameters were computed using structure-from-motion (SfM) (Schönberger and Frahm, 2016) based on the static scene. Subsequently, we randomly posed the robot and used only the color sensor to capture images from the calculated view angles. We restricted the joint angles in $\frac{\pi}{6},\frac{\pi}{6}]$ to capture 100 images, then posed the robot without joint limitations to capture additional 1,000 images. We employed the segment anything model (SAM) (Kirillov et al., 2023) to generate masks and segment the robot from the background.
Table 2: Comparison of Chamfer distances without joint limitations.
Method Configuration $\theta^{(a)}$ Configuration $\theta^{(c)}$ Configuration $\theta^{(d)}$ Configuration $\theta^{(e)}$
Ours $3.6\times 10^{-4}$ $4.4\times 10^{-3}$ $4.5\times 10^{-4}$ $3.4\times 10^{-4}$
Schulze and Lipson (2024) $1.7\times 10^{-2}$ $5.3\times 10^{-2}$ $1.3\times 10^{-2}$ $1.3\times 10^{-2}$
Chen et al. (2022) $4.3\times 10^{-4}$ $5.8\times 10^{-4}$ $2.3\times 10^{-3}$ $1.7\times 10^{-3}$
@@
The results of physical experiments are shown in Figure 9. We randomly selected three sets of joint angles that were not in the training data, rendered the corresponding robot images using the trained model, and compared them to the ground truth. The comparison reveals that the proposed method accurately represents the robot’s morphology and surface color, further validating the effectiveness of our approach.
Conducting self-modeling with real-world data is more challenging than with simulation data. Thus, the real-world results are not as accurate as those obtained from simulations. This can be attributed to several factors: (1) Potential inaccuracies in the camera’s intrinsic and extrinsic parameters calculated by SfM, (2) Imperfect segmentation of the robots from the background in certain images, and (3) A smaller volume of data is available in real-world experiments compared to simulation. A failure example is illustrated in Figure 10. In the presented example, the robot’s end effector was observed to be detached from its main body. Furthermore, the morphology of the robot’s end-effector was also found to be inaccurate. Possible reasons for this failure include the limited quantity of the images we collected. Another potential reason is our assumption that each link has a similar size, whereas the volumes of the robot’s links differ significantly in reality.
Figure 10: A failure example of the physical experiment.
@@ 4.3 Applications
In this section, we use the proposed self-model to complete some downstream tasks: reaching a target, touching a target with the robot’s surface, and reaching a target while avoiding an obstacle. We also show that our self-model can be used to solve the inverse kinematics problem.
@@ 4.3.1 Reaching a target
A green ball was placed within the robot’s workspace, requiring the robot to use its self-model to determine how to make contact with the ball using a specific part of its body. Our neural bone representation simplifies identifying specific parts of the robot and enables targeted control to reach the desired object. We use the robot’s end effector as an example. Starting from a resting configuration, the joint angles are progressively adjusted to minimize the distance between the target and the center of the neural bone associated with the end effector. This is accomplished by fixing the self-model parameters and applying gradient descent to automatically optimize the joint angles. In practice, the AdamW (Loshchilov and Hutter, 2019) optimizer is utilized, and the loss function is defined as the mean squared error (MSE) between the bone center and the target position:
[…]
where $p_{b}$ is the position of the center of the specified bone, and $p_{t}$ is the position of the target.
In order to keep joint angles within the range limit, we use a sine function to constrain joint values within [1,1][-1,1] and avoid introducing discontinuities. We use a learning rate of 0.01 to avoid substantial changes in the optimizable joint values. Then we scale it to the actual range limit. Unlike the method in Chen et al. (2022), by altering the neural bone applied in the loss function, the robot can be controlled to reach the target using different parts without the need for additional training.
4.3.2 Touching a target with surface
The above method is only applicable when the target is a point in space rather than an actual sphere, as the target position will end up inside the robot link. Therefore, a different approach is required if the target is an actual sphere. We regard all 3D Gaussians as point clouds. Instead of computing the distance between the center of the target sphere and the center of the neural bone, we compute the distance between the surface of the target sphere and the points belonging to the specific link and minimize it. Again, we use the end effector as an example. We use the skinning weights derived by (13) of each 3D Gaussian to determine which part they belong to. This process is formulated as
$\mathcal{M}_{i}=\arg\max_{m\in{1,2,...,M}}(w_{im})$ (27)
where $w_{i}$ is the skinning weights of 3D Gaussian $i$ with respect to all neural bones, and $\mathcal{M}_{i}$ is the index of the part that 3D Gaussian $i$ belongs to. Then, the loss function is
[…]
where $\mathcal{M}_{ee}$ is the index of the end effector, $p_{target}$ is the position of the target and $r$ is radius of the target sphere. We stop optimization to avoid collision with the target when the minimum distance is less than 0.001.
Figure 11: Results of applications. a Snapshots and joint angles of the robot with respect to step number. b Snapshots from different perspectives, joint angles of the robot, and the minimum distance between the target and the end effector of the robot with respect to step number. c Snapshots from different perspectives, joint angles of the robot, and the minimum distance between the obstacle and the robot with respect to step number. The unit of joint angles is radians. At the beginning, all joints of the robot are set to 0. Using solely the self-model, the robot is able to find joint configurations that can fulfill the tasks. d Results of inverse kinematics estimation. The ground truth joint angles are [-0.204, -0.9319, -0.3448, -1.5012, 0.4474, 0.105, 0.3166]. The estimated joint angles are [-0.2127, -0.9321, -0.3051, -1.5050, 0.4863, 0.0258, 0.3113]. Joint angles in radians are slightly different, but the morphology of the robot is close to the ground truth.
@@ 4.3.3 Reaching a target while avoiding an obstacle
An obstacle is placed at a location the robot would pass through after completing the first task of reaching the target. The robot is required to find a new joint configuration to reach the target, as the original configuration would result in a collision with the obstacle. To address this, an additional loss function is introduced alongside the MSE loss function used in the previous task. All 3D Gaussians are treated as point clouds, and the distances $D$ between the points and the obstacle are computed. The loss function is then formulated as
@@
where $\varepsilon$ is a threshold that controls how close the robot can get to the obstacle. In our experiment, about 50,000 3D Gaussians were used in the self-model. All 3D Gaussians were forced to keep a distance from the obstacle. When faced with more challenging situations, such as the presence of multiple obstacles in the space, our proposed simplified method may fail and can not generate a collision-free trajectory. However, since the self-model can compute the distance from each obstacle to each 3D Gaussian ellipsoid, we can modify mature motion planning algorithms to generate reliable, collision-free trajectories in complex situations.
The above three applications are demonstrated in Figure 11. Images in the first row show how our self-model controls the end effector of the robot to reach the target, shown here as a green ball. The robot’s final pose in this task is shown in the fifth image of the first row (Movie S1). In the second and third lines, we illustrate how the robot is controlled to touch the target sphere with its surface (Movie S2 and Movie S3). The joint angles and minimum distance between the 3D Gaussians and the target surface are shown in the last column. Next, an obstacle was placed at the position of the robot’s fourth link in the final pose of task 1, and images in the fourth and fifth rows show two perspectives of our self-model guiding the robot to complete the target-reaching task while avoiding the obstacle. Clearly, the robot adopts a new pose to avoid the obstacle and reach the target (Movie S4 and Movie S5). The minimum distance between the obstacle and the robot’s 3D Gaussian representations is also shown in the last column of Figure 11(c).
Compared with traditional methods, our method achieves control of unknown robots to accomplish tasks. A traditional method needs to manually create both a kinematic and a geometric model of the unknown robot before the implementation of the control algorithm. However, our method can autonomously execute the complete pipeline from modeling to control.
@@ 4.3.4 Robot inverse kinematics estimation
With our self-model, we can estimate the robot’s inverse kinematics. Given images of the robot, our self-model can predict a joint configuration to pose the robot similarly. After freezing self-model parameters and treating joint angles as optimizable, we employ gradient descent to approximate the joint configuration. The optimization process begins from the rest pose and terminates when the magnitude of change in joint angles falls below a predefined threshold. Comparing rendered and ground truth images, we optimize joint angles to achieve reasonable values. In practice, we use two images to estimate the robot’s pose. Our estimation results are illustrated in Figure 11(d). Although estimated joint angles differ slightly from the actual angles, rendered images at the estimated joint angles match ground-truth images. Our self-model effectively poses the robot in alignment with the ground truth. Additional images from varied view angles could further improve joint angle estimation accuracy.
@@ 5 Discussion
Our method achieves robot self-modeling and can be used to control the robot. However, the method can be further improved in several ways. (1) Our method is based on the assumption of rigid links of robots, which limits its direct applicability to soft and continuous robots. Extending this method to the soft robot is an important future work. (2) A better initialization of our neural bones is possible. The current method requires a predefined number of bones to initialize the structure. This number is an important hyperparameter that affects the training process. We also assume that each bone is similar in size, but in reality, the size of robot links may vary, necessitating multiple bones to represent a single link. (3) Our method may not always accurately reconstruct the robot at certain joint configurations, especially when the joint’s movement range is unrestricted. Besides, due to the limitations imposed on joint angles, the range of applications for the models built using our approach is restricted, as some tasks would unavoidably require the robot to move out of the joint constraint. It is worth researching how to obtain a more accurate self-model across all joint configurations. (4) It is important to find a method to map the 3DGS self-model to a mesh model and modify it for compatibility with common robot control pipelines, such as MOVEIT (Coleman et al., 2014) or current reinforcement learning toolkits. (5) We need multiple views to reconstruct the model, and the number of views, along with the camera parameters, can affect the final result. Masks of robots are also required. Researching how to self-model the robot with high quality using fewer views and without known camera parameters and masks is a valuable direction to explore. (6) Several other hyperparameters are required to train the self-model, for example, the weights of the loss components. Limiting the number of hyperparameters is another task for the future.
6 Conclusion
We propose a 3DGS-based robot self-modeling method. This approach achieves high-quality modeling of the robot’s morphology, surface color, and kinematics. Compared to the previous method that requires RGB images only, our method produces results with much better geometric accuracy. Our method initiates modeling from the robot’s rest pose. We employ learnable neural ellipsoid bones of similar volume to represent the underlying articulated structure and kinematics link of the robot. A kinematic network is trained to capture the robot’s kinematic properties. Each 3D Gaussian’s movement is governed by transformations of neural bones, enabling full model control. Rendering 3D Gaussians into images enables us to visualize the robot’s morphology and surface color. Our self-model accurately reconstructs the robot across various joint configurations, achieving high-quality mesh extractions and image synthesis.
The self-model developed using our method can be applied in downstream tasks. With this self-model, the robot is capable of solving motion-planning tasks. Our neural bone representation directly models the robot’s structure, allowing us to control different parts of the robot by specifying which bone to optimize in the loss function without requiring additional information or training to adjust the self-model. Since our self-model can synthesize high-quality images, it can also estimate the robot’s inverse kinematics by rendering and comparing with ground truth images.
{dci}
@@ 6 Conclusion
The author(s) declared no potential conflicts of interest with respect to the research, authorship, and/or publication of this article.
@@
Movie S2: The robot is controlled by its self-model to touch the green target sphere with its surface at viewpoint 1.
Movie S3: The robot is controlled by its self-model to touch the green target sphere with its surface at viewpoint 2.
Movie S4: The robot is controlled by its self-model to reach the green target while avoiding the red obstacle at viewpoint 1.
Movie S5: The robot is controlled by its self-model to reach the green target while avoiding the red obstacle at viewpoint 2.